# Exact Bayesian Inference for Markov Switching Diffusions

### v1

@@ 3.2 Hidden Data Update
$\begin{split}&\frac{\alpha_{(V_{R},Z,Y)}(\{v_{r^{\dagger}}^{\dagger},z^{\dagger},y^{\dagger}\},\{v_{r},z,y\})}{\alpha_{(V_{R},Z,Y)}(\{v_{r},z,y\},\{v_{r^{\dagger}}^{\dagger},z^{\dagger},y^{\dagger}\})}\\ &=\frac{\kappa(v_{r},z,y|v_{s})}{\kappa(v_{r^{\dagger}}^{\dagger},z^{\dagger},y^{\dagger}|v_{s})}\frac{\pi(v_{r^{\dagger}}^{\dagger},z^{\dagger},y^{\dagger}|v_{s},\theta,\lambda)}{\pi(v_{r},z,y|v_{s},\theta,\lambda)}\\ &=\frac{\kappa(v_{r}|v_{s},y)}{\kappa(v_{r^{\dagger}}^{\dagger}|v_{s},y^{\dagger})}\frac{\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau^{\dagger}}\pi(z_{(\dot{\tau},\ddot{\tau})}^{\dagger},v_{\ddot{\tau}}^{\dagger}|v_{\dot{\tau}}^{\dagger},y_{\dot{\tau}}^{\dagger},\theta)}{\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}\pi(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta)}\\ &=\frac{\overbrace{\kappa(v_{r}|v_{s},y)\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau^{\dagger}}d_{\theta}(z_{(\dot{\tau},\ddot{\tau})}^{\dagger},v_{\{\dot{\tau},\ddot{\tau}\}}^{\dagger},y_{\dot{\tau}}^{\dagger})}^{c_{1}}}{\underbrace{\kappa(v_{r^{\dagger}}^{\dagger}|v_{s},y^{\dagger})\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}d_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})}_{c_{2}}}\frac{\overbrace{\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau^{\dagger}}q_{\theta}(z_{(\dot{\tau},\ddot{\tau})}^{\dagger},v_{\{\dot{\tau},\ddot{\tau}\}}^{\dagger},y_{\dot{\tau}}^{\dagger})}^{p_{1}}}{\underbrace{\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}q_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})}_{p_{2}}},\end{split}$ (29)
@@
which gives a valid 2-coin factorization. Notice that we can generate a coin with probability $p_{2}=\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}q_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$ by flipping subordinate $q_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$-coins with the Poisson coin algorithm, and returning heads if all subordinate coin flips are heads. Conversely, we can return tails as soon as one of the subordinate coin flips is tails. Moreover, we can split $\{v_{r^{\dagger}}^{\dagger},z^{\dagger},y^{\dagger}\}$ into separate sections at $\{\dot{s}\in s:y_{\dot{s}}=y_{\dot{s}}^{\dagger}\}$, and accept or reject those separately. We expand on that in the localized version given in Supplement B.
@@ 3.3 Diffusion Parameter Update
We carry out the update to $\pi(\theta|v_{\tau},z)$ by way of a Barker-within-Gibbs step with generic proposal density $\kappa(\theta^{\dagger}|\theta)$. Such a proposal may be adjusted adaptively, e.g. by way of Adapting Increasingly Rarely (chimisov2018air) or other adaptive MCMC methods. The proposed value $\theta^{\dagger}$ has acceptance odds
@@
$\begin{split}\frac{\alpha_{\Theta}(\theta,\theta^{\dagger})}{\alpha_{\Theta}(\theta^{\dagger},\theta)}&=\frac{\kappa(\theta|\theta^{\dagger})}{\kappa(\theta^{\dagger}|\theta)}\frac{\pi(\theta^{\dagger}|v_{\tau},z,y)}{\pi(\theta|v_{\tau},z,y)}\\ &=\frac{\kappa(\theta|\theta^{\dagger})}{\kappa(\theta^{\dagger}|\theta)}\frac{\pi(\theta^{\dagger})}{\pi(\theta)}\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}\frac{\pi(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta^{\dagger})}{\pi(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta)},\end{split}$ (30)
@@
where the immediate way of constructing a 2-coin algorithm would be to factorize as in the hidden data update, such that $\pi(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta^{\dagger})=d_{\theta}(z_{(\dot{\tau},\ddot{\tau})},v_{\{\dot{\tau},\ddot{\tau}\}},y_{\dot{\tau}})q_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$. This solution has two major downsides, the first of is that the integrand in
@@
$\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}q_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}\exp\left\{\int_{\dot{\tau}}^{\ddot{\tau}}{\tilde{\varphi}}^{\downarrow}_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})-\tilde{\varphi}_{\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\mathop{}\!\mathrm{d}t\right\}$ (31)
@@
accumulates linearly in $\omega$. Therefore, the compount coin has probability of heads of order $\mathcal{O}(e^{-\omega})$, and following (18), the superordinate 2-coin algorithm has expected number of iterations of order $\mathcal{O}(e^{\omega})$. Secondly, the cost of this 2-coin algorithm does not decrease in the step size $|\theta^{\dagger}-\theta|$. This issue is partially due to the fact that the locality of the $\theta$-update is not reflected in the factorization.
@@
A partial solution consists of considering ratios of intractable quantities, such as in the exchange algorithm of murray2012mcmc. If we define
$\xi_{t}=\tilde{\varphi}_{\theta^{\dagger}}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})-\tilde{\varphi}_{\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\qquad(t\in(\dot{\tau},\ddot{\tau}))$ (32)
with positive and negative parts $\xi_{t}^{(+)}$ and $\xi_{t}^{(-)}$, we obtain the valid 2-coin factorization
@@
$\begin{split}\frac{\alpha_{\Theta}(\theta,\theta^{\dagger})}{\alpha_{\Theta}(\theta^{\dagger},\theta)}&=\frac{\kappa(\theta|\theta^{\dagger})}{\kappa(\theta^{\dagger}|\theta)}\frac{\pi(\theta^{\dagger})}{\pi(\theta)}\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}\frac{h_{\theta^{\dagger}}(v_{\{\dot{\tau},\ddot{\tau}\}},y_{\dot{\tau}})}{h_{\theta}(y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})}\exp\left\{-\int_{\dot{\tau}}^{\ddot{\tau}}\xi_{t}\mathop{}\!\mathrm{d}t\right\}\\ &=\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}\frac{\overbrace{\{\kappa(\theta|\theta^{\dagger})\pi(\theta^{\dagger})\}^{1/(|\tau|-1)}h_{\theta^{\dagger}}(v_{\{\dot{\tau},\ddot{\tau}\}},y_{\dot{\tau}})}^{c_{1}^{(\dot{\tau},\ddot{\tau})}}}{\underbrace{\{\kappa(\theta^{\dagger}|\theta)\pi(\theta)\}^{1/(|\tau|-1)}h_{\theta}(y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})}_{c_{2}^{(\dot{\tau},\ddot{\tau})}}}\frac{\overbrace{e^{-\int_{\dot{\tau}}^{\ddot{\tau}}\xi_{t}^{(+)}\mathop{}\!\mathrm{d}t}}^{p_{1}^{(\dot{\tau},\ddot{\tau})}}}{\underbrace{e^{-\int_{\dot{\tau}}^{\ddot{\tau}}\xi_{t}^{(-)}\mathop{}\!\mathrm{d}t}}_{p_{2}^{(\dot{\tau},\ddot{\tau})}}}.\end{split}$ (33)
@@
Loose bounds on the integrands in the range $t\in(\dot{\tau},\ddot{\tau})$ are given by
$\xi_{t}^{(+)}\leq{\tilde{\varphi}}^{\uparrow}_{\theta^{\dagger}}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})-{\tilde{\varphi}}^{\downarrow}_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}}),$ (34)
$\xi_{t}^{(-)}\leq{\tilde{\varphi}}^{\uparrow}_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})-{\tilde{\varphi}}^{\downarrow}_{\theta^{\dagger}}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}}).$ (35)
@@
Let $\pi(\theta)$ be supported on a compact set $\mathcal{T}\subset\mathbf{R}$ with $|\partial_{\theta}\tilde{\varphi}_{\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})|$ uniformly bounded, fix the observation interval $\ddot{s}-\dot{s}$ for all $(\ddot{s}\sim\dot{s})\in s$, and suppose that $\pi(\theta|v_{s})$ satisfies a Bernstein-von-Mises theorem as $|s|,\omega\to\infty$, such that for the proposal $\theta^{\dagger}|\theta\sim\operatorname{\mathrm{Unif}}\left[\theta\pm\varsigma\right]$, the optimal step size $\varsigma$ is of order $\mathcal{O}(1/\sqrt{\omega})$. Then, $\int_{0}^{\omega}|\xi_{t}|\mathop{}\!\mathrm{d}t=\mathcal{O}(\sqrt{\omega})$.
@@
$|\xi_{t}|\leq|\theta^{\dagger}-\theta|\sup_{\theta\in\mathcal{T}}\left\lvert\partial_{\theta}\tilde{\varphi}_{\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\right\rvert,\qquad(t\in(\dot{\tau},\ddot{\tau}))$ (36)
and since $\theta^{\dagger}-\theta\sim\operatorname{\mathrm{Unif}}\left[0,\varsigma\right]$, $|\theta^{\dagger}-\theta|\leq\varsigma=\mathcal{O}(1/\sqrt{\omega})$. By the uniform bound on the gradient, the claim follows. ∎
While the bounded gradient assumption is stringent, a similar result could be shown to hold on average, and it illustrates that the alternative factorization exploits posterior concentration. Notice that such a factorization could also be adopted in the hidden data update - this would be particularly useful if that update is highly localized, such that the integrands contributed by state and proposal cancel out. We further improve the scaling of the parameter update by applying the *divide-and-conquer Bernoulli factory* of stumpf2024bf, which takes as inputs the full collection of weights $\{c_{1}^{(\dot{\tau},\ddot{\tau})}\}$ and $\{c_{2}^{(\dot{\tau},\ddot{\tau})}\}$ as well as coins $\{p_{1}^{(\dot{\tau},\ddot{\tau})}\}$ and $\{p_{2}^{(\dot{\tau},\ddot{\tau})}\}$, and constructs the coin hierarchically from smaller batches. In stumpf2024bf, it was observed that the combination of those remedies improves the cost of the $\theta$-update to as low as $\mathcal{O}(\omega\log\omega)$ for simple diffusion models, which we expect to hold for switching models more broadly.

### v2

@@ 3.2 Hidden Data Update
$\begin{split}&\frac{\alpha_{(V_{R},Z,Y)}(\{v_{r^{\dagger}}^{\dagger},z^{\dagger},y^{\dagger}\},\{v_{r},z,y\})}{\alpha_{(V_{R},Z,Y)}(\{v_{r},z,y\},\{v_{r^{\dagger}}^{\dagger},z^{\dagger},y^{\dagger}\})}\\ &=\frac{\kappa(v_{r},z,y|v_{s})}{\kappa(v_{r^{\dagger}}^{\dagger},z^{\dagger},y^{\dagger}|v_{s})}\frac{\pi(v_{r^{\dagger}}^{\dagger},z^{\dagger},y^{\dagger}|v_{s},\theta,\lambda)}{\pi(v_{r},z,y|v_{s},\theta,\lambda)}\\ &=\frac{\kappa(v_{r}|v_{s},y)}{\kappa(v_{r^{\dagger}}^{\dagger}|v_{s},y^{\dagger})}\frac{\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau^{\dagger}}\pi(z_{(\dot{\tau},\ddot{\tau})}^{\dagger},v_{\ddot{\tau}}^{\dagger}|v_{\dot{\tau}}^{\dagger},y_{\dot{\tau}}^{\dagger},\theta)}{\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}\pi(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta)}\\ &=\frac{\overbrace{\kappa(v_{r}|v_{s},y)\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau^{\dagger}}d_{\theta}(z_{(\dot{\tau},\ddot{\tau})}^{\dagger},y_{\dot{\tau}}^{\dagger},v_{\{\dot{\tau},\ddot{\tau}\}}^{\dagger})}^{c_{1}}}{\underbrace{\kappa(v_{r^{\dagger}}^{\dagger}|v_{s},y^{\dagger})\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}d_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})}_{c_{2}}}\frac{\overbrace{\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau^{\dagger}}q_{\theta}(z_{(\dot{\tau},\ddot{\tau})}^{\dagger},y_{\dot{\tau}}^{\dagger},v_{\{\dot{\tau},\ddot{\tau}\}}^{\dagger})}^{p_{1}}}{\underbrace{\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}q_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})}_{p_{2}}},\end{split}$ (29)
@@
which gives a valid 2-coin factorization. Notice that we can generate a coin with probability $p_{2}=\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}q_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$ by flipping subordinate $q_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$-coins with the Poisson coin algorithm, and returning 1 if all subordinate coin flips are 1. Conversely, we can return 0 as soon as one of the subordinate coin flips is 0. Moreover, we can split $\{v_{r^{\dagger}}^{\dagger},z^{\dagger},y^{\dagger}\}$ into separate sections at $\{\dot{s}\in s:y_{\dot{s}}=y_{\dot{s}}^{\dagger}\}$, and accept or reject those separately. We expand on that in the localized version given in Supplement B.
@@ 3.3 Diffusion Parameter Update
We carry out the update to $\pi(\theta|v_{\tau},z,y)$ by way of a Barker-within-Gibbs step with generic proposal density $\kappa(\theta^{\dagger}|\theta)$. Such a proposal may be adjusted adaptively, e.g. by way of Adapting Increasingly Rarely [11] or other adaptive MCMC methods. The proposed value $\theta^{\dagger}$ has acceptance odds
@@
$\begin{split}\frac{\alpha_{\Theta}(\theta,\theta^{\dagger})}{\alpha_{\Theta}(\theta^{\dagger},\theta)}&=\frac{\kappa(\theta|\theta^{\dagger})}{\kappa(\theta^{\dagger}|\theta)}\frac{\pi(\theta^{\dagger}|v_{\tau},z,y)}{\pi(\theta|v_{\tau},z,y)}=\frac{\kappa(\theta|\theta^{\dagger})}{\kappa(\theta^{\dagger}|\theta)}\frac{\pi(\theta^{\dagger})}{\pi(\theta)}\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}\frac{\pi(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta^{\dagger})}{\pi(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta)},\end{split}$ (30)
@@
where the immediate way of constructing a 2-coin algorithm would be to factorize as in the hidden data update, such that $\pi(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta^{\dagger})=(d_{\theta}\times q_{\theta})(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$. This solution has two major downsides, the first of which is that the integrand in
@@
$\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}q_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=\exp\left\{\sum_{(\dot{\tau}\sim\ddot{\tau})\in\tau}\int_{\dot{\tau}}^{\ddot{\tau}}{\tilde{\varphi}}^{\downarrow}_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})-\tilde{\varphi}_{\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\mathop{}\!\mathrm{d}t\right\}$ (31)
@@
accumulates linearly in $\omega$. Therefore, the compound coin has probability of heads of order $e^{-\mathcal{O}(\omega)}$, and following (18), the superordinate 2-coin algorithm has expected number of iterations of order $e^{\mathcal{O}(\omega)}$. Secondly, the cost of this 2-coin algorithm does not decrease in the step size $|\theta^{\dagger}-\theta|$. This is due to the fact that the locality of the $\theta$-update is not reflected in the factorization.
@@
A partial solution consists of working with ratios of intractable quantities, as in the exchange algorithm of [37]. If we define
$\xi_{\theta^{\dagger},\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=(\tilde{\varphi}_{\theta^{\dagger}}-\tilde{\varphi}_{\theta})(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\qquad(t\in(\dot{\tau},\ddot{\tau}))$ (32)
we obtain the valid 2-coin factorization
@@
$\begin{split}\frac{\alpha_{\Theta}(\theta,\theta^{\dagger})}{\alpha_{\Theta}(\theta^{\dagger},\theta)}&=\frac{\kappa(\theta|\theta^{\dagger})}{\kappa(\theta^{\dagger}|\theta)}\frac{\pi(\theta^{\dagger})}{\pi(\theta)}\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}\frac{h_{\theta^{\dagger}}(y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})}{h_{\theta}(y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})}\exp\left\{-\int_{\dot{\tau}}^{\ddot{\tau}}\xi_{\theta^{\dagger},\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\mathop{}\!\mathrm{d}t\right\}\\ &=\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}\frac{\overbrace{\{\kappa(\theta|\theta^{\dagger})\pi(\theta^{\dagger})\}^{1/(|\tau|-1)}h_{\theta^{\dagger}}(y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})}^{c_{1}^{(\dot{\tau},\ddot{\tau})}}}{\underbrace{\{\kappa(\theta^{\dagger}|\theta)\pi(\theta)\}^{1/(|\tau|-1)}h_{\theta}(y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})}_{c_{2}^{(\dot{\tau},\ddot{\tau})}}}\frac{\overbrace{e^{-\int_{\dot{\tau}}^{\ddot{\tau}}\xi_{\theta^{\dagger},\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})^{(+)}\mathop{}\!\mathrm{d}t}}^{p_{1}^{(\dot{\tau},\ddot{\tau})}}}{\underbrace{e^{-\int_{\dot{\tau}}^{\ddot{\tau}}\xi_{\theta,\theta^{\dagger}}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})^{(+)}\mathop{}\!\mathrm{d}t}}_{p_{2}^{(\dot{\tau},\ddot{\tau})}}},\end{split}$ (33)
@@
where the superscript $(+)$ denotes the positive part. Loose bounds on the integrands in the range $t\in(\dot{\tau},\ddot{\tau})$ are given by
$\xi_{\theta^{\dagger},\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\leq({\tilde{\varphi}}^{\uparrow}_{\theta^{\dagger}}-{\tilde{\varphi}}^{\downarrow}_{\theta})(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}}).$ (34)
@@
Let $\pi(\theta)$ be supported on a compact set $\mathcal{T}\subset\mathbf{R}$ with $|\partial_{\theta}\tilde{\varphi}_{\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})|$ uniformly bounded, fix the observation interval $\ddot{s}-\dot{s}$ for all $(\ddot{s}\sim\dot{s})\in s$, and set the proposal $\theta^{\dagger}|\theta\sim\operatorname{\mathrm{Unif}}\left[\theta\pm\varsigma/\sqrt{\omega}\right]$. Then, $\int_{0}^{\omega}|\xi_{\theta^{\dagger},\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})|\mathop{}\!\mathrm{d}t=\mathcal{O}(\sqrt{\omega})$.
@@
$|\xi_{\theta^{\dagger},\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})|\leq|\theta^{\dagger}-\theta|\sup_{\theta\in\mathcal{T}}\left\lvert\partial_{\theta}\tilde{\varphi}_{\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\right\rvert,\qquad(t\in(\dot{\tau},\ddot{\tau}))$ (35)
and since $\theta^{\dagger}-\theta\sim\operatorname{\mathrm{Unif}}\left[0,\varsigma/\sqrt{\omega}\right]$, $|\theta^{\dagger}-\theta|\leq\varsigma/\sqrt{\omega}$. By the uniform bound on the gradient, the claim follows. ∎
Scaling the step size as $\mathcal{O}(\omega^{-1/2})$ is justified by the typical Bernstein-von Mises posterior concentration rate of $\omega^{-1/2}$ when appending observations at constant intervals, and usually results in an acceptance rate of constant order as $\omega\to\infty$ [41]. The bounded gradient assumption is more stringent, but a similar result could be shown to hold on average, and it illustrates that the alternative factorization exploits posterior concentration. Such a factorization could also be adopted in the hidden data update - this would be particularly useful if that update is highly localized, such that the integrands contributed by state and proposal cancel out. We further improve the scaling of the parameter update by applying the divide-and-conquer Bernoulli factory of [44], which takes as inputs the full collection of weights $\{c_{1}^{(\dot{\tau},\ddot{\tau})}\}$ and $\{c_{2}^{(\dot{\tau},\ddot{\tau})}\}$ as well as coins $\{p_{1}^{(\dot{\tau},\ddot{\tau})}\}$ and $\{p_{2}^{(\dot{\tau},\ddot{\tau})}\}$, and constructs the acceptance coin hierarchically from smaller batches. In [44], it was observed that the combination of those remedies improves the cost of the $\theta$-update to as low as $\mathcal{O}(\omega)$ for ordinary diffusion models, which we expect to hold for switching models more broadly, as opposed to a scaling of $e^{\mathcal{O}(\sqrt{\omega})}$ for the naive 2-coin implementation.